OpenStudy (anonymous):

Find the length of the missing side of this right triangle. @heather040200

OpenStudy (anonymous):



OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (dumbcow):

sounds like Pythagoras may be involved here...

OpenStudy (anonymous):

4^2+7^2=c^2

OpenStudy (anonymous):

8"